Need a solution for breaking out two ticket prices. $49 for a child ticket and $55 for an adult ticket. There are only those two ticket prices and we need a count of how many child and adult tickets there are.
Total tickets are $56,290
Is there any formulae that could do this?
Thank you!

Answer by richwmiller(17219)

49c+55a=56290
a and c both must be positive integers (whole numbers)
if a=0 then c = 1148.8
if c=0 then a = 1023.5
so a < 1023.5 and c < 1148.8
a = 49n+39 and c = 1105-55n
a = 88, c = 1050, n=1
a = 137, c = 995, n = 2
a = 186, c = 940, n = 3
a = 235, c = 885, n = 4
a = 284, c = 830, n = 5
we could pin it down to one answer if we knew the total number of tickets sold.